### v2

# Learning policies for resource allocation in business processes

Abstract
Efficient allocation of resources to activities is pivotal in executing business processes but remains challenging. While resource allocation methodologies are well-established in domains like manufacturing, their application within business process management remains limited. Existing methods often do not scale well to large processes with numerous activities or optimize across multiple cases. This paper aims to address this gap by proposing two learning-based methods for resource allocation in business processes. The first method leverages Deep Reinforcement Learning (DRL) to learn near-optimal policies by taking action in the business process. The second method is a score-based value function approximation approach, which learns the weights of a set of curated features to prioritize resource assignments. To evaluate the proposed approaches, we first designed six distinct business processes with archetypal process flows and characteristics. These business processes were then connected to form three realistically sized business processes. We benchmarked our methods against traditional heuristics and existing resource allocation methods. The results show that our methods learn adaptive resource allocation policies that outperform or are competitive with the benchmarks in five out of six individual business processes. The DRL approach outperforms all benchmarks in all three composite business processes and finds a policy that is, on average, 13.1% better than the best-performing benchmark.
keywords:
Resource allocation, business process optimization, deep reinforcement learning, Bayesian optimization
[…]
1 Introduction
Efficiently allocating resources to activities that must be executed in a business process at a particular moment at run-time is crucial for organizations to achieve operational outcomes. Optimally utilizing resources to their full capacity maximizes their productivity and enhances an organization’s profitability and competitiveness.
In recent years, the main contribution to data-driven optimization of business processes has been seen in Prescriptive Processing Monitoring (PrPM) kubrak_prescriptive_2021 . This research area has seen a strong increase in attention and aims to make recommendations for a case (i.e., an instance of the process) that lead to the best outcome. For example, several PrPM methods recommend the best resource to perform a specific activity for running cases to minimize the case’s completion time sindhgatta_context-aware_2016 ; wibisono_–fly_2015 or cycle time thomas_online_2017 . However, the main limitation of PrPM is that it only optimizes for individual cases and does not consider the effect of intervention for one case on other cases kubrak_prescriptive_2021 .
Data-driven business process optimization (BPO) methods aim to optimize the performance of the complete process and not just individual cases (e.g.,  park_prediction-based_2019 ; huang_reinforcement_2011 ; firouzian_cycle_2019 ; zbikowski_deep_2023 ). In BPO, operational decisions are made to optimize process key performance indicator(s) of interest, such as the mean cycle time or throughput, instead of single case outcome. Take an example of a business process given in Fig. 1. This process has two activities, Accept Application (AA) and Reject Application (RA), and two resources, $r_{1}$ and $r_{2}$. Activity AA can only be performed by resource $r_{1}$, and the activity RA can be performed by both resources. In a situation where two cases are present, case A at activity AA and case B at activity RA, a PrPM method would recommend using resource $r_{1}$ for both cases as it the most efficient resource. However, assigning resource $r_{1}$ to case B means that case A needs to wait until resource $r_{1}$ becomes available again. From a process perspective, assigning resource $r_{2}$ to case B and resource $r_{1}$ to case A would be efficient as both cases can be processed simultaneously and are not competing with each other over the available resources.
Figure 1: An example of a business process model of a loan application process with two activities and two resources. Cases arrive according to some distribution with rate $\lambda$, and the loans are either accepted (AA) or rejected (RA). Each activity can be performed by both resources $r_{1}$ and $r_{2}$, according to a processing time distribution with mean $\mathbb{E}(X_{i})$.
The online resource allocation problem in business processes is a dynamic task assignment problem spivey_dynamic_2004 . Several approaches have been proposed to solve this problem. Park and Song park_prediction-based_2019 predicted the next activities and processing times of activity-resource pairs in running cases and solved a min-cost max-flow network problem to find optimal resource assignments. Huang et al. huang_reinforcement_2011 used Reinforcement Learning (RL) to learn the effective resource assignments. This approach was extended by Firouzian et al. firouzian_cycle_2019 by including entropy-based features mined from the event log. However, the tabular Q-learning method used in these papers is unsuitable for problems with large state spaces, such as business processes containing numerous activities and resources. Žbiwkowski et al. zbikowski_deep_2023 partly solved this problem using Deep Reinforcement Learning (DRL), where a neural network replaces the Q-table to maximize the number of completed cases. Their method, however, learns feasible actions through experience, which becomes computationally expensive in larger business processes and may not converge to competitive policies.
In this paper, we present two learning methods for resource allocation in business processes to minimize the mean cycle time of cases. The first method leverages DRL to learn the goodness of assignments by interacting with the simulator of a business process. It can autonomously derive high-dimensional features directly from the state information. The second method is a score-based value function approximation (SVFA) method, which leverages domain expertise to curate process features and learns weights for these features to rank different resource assignments.
To evaluate our models, we first created six distinct business process models with archetypal business process flows and characteristics, such as parallel and choice patterns, varying resource utilization rates, and resource competition. Subsequently, these six business processes were combined in sequential and parallel order to create three realistically-sized business processes. We benchmarked our methods with traditional heuristics and existing BPO methods. The results show that our learning methods outperform or are competitive with the benchmarks in five out of the six business processes. Our DRL method outperforms all of the benchmarks in all three of the composite business processes.
The outline of this paper is as follows. Section 2 gives an overview of the related work. Section 3 introduces definitions and concepts related to the execution of business process models. Section 4 presents our two learning methods to solve the resource allocation problem. Section 5 introduces the business processes used for evaluation and details the training and evaluation protocol. Section 6 concludes our work.
2 Related work
In business process management, much research has been done on the behavior and allocation of resources. Several methods study the behavior of resources resulting from the current resource allocation policy. This behavior can be evaluated using a set of resource behavior indicators, such as resource preference and productivity pika_extensible_2014 , or from different perspectives huang_resource_2012 . López-Pintado et al. lopez-pintado_prosimos_2023 developed specialized tools for discovering resource profiles to differentiate resource behavior within roles. Furthermore, the current resource allocation policy may also be used to assess the efficiency of a business process duran_analysis_2019 .
While these methods generate valuable insights into the current resource allocation policy, they do not consider optimizing existing policies. Our work focuses on improving the resource allocation policy itself, and therefore, we will discuss studies that propose algorithms for resource allocation. We can distinguish between rule-based and dynamic resource allocation methods in the literature.
Several rule-based methods use process mining techniques to determine static resource allocation strategies. Xu et al. xu_resource_2008 assigned resources to activities based on the workflow of the process model. Other methods use resource-related information mined from the event log in various dimensions to create a ranking of resource assignments, which is used to assign resources to activities at run-time zhao_optimization_2015 ; arias_framework_2016 ; kuchar_automatic_2016 . A generic genetic algorithm was also used for a resource assignment scheme for a specific process based on colored Petri nets si_petri_2018 . These methods provide a generalized resource allocation scheme for the business process model but are not adaptive to the specific circumstances of the process, such as the number of available tasks and resources.
Dynamic methods make assignments based on the current circumstances of the process. Park and Song park_prediction-based_2019 proposed a scheduling approach that uses an offline prediction model to predict the next activities and processing times of resource assignments of a running case. These predictions serve as input for the minimum cost and maximum flow algorithm, which finds optimal resource-activity pairs at each decision step. However, the method only optimizes for the current decision step and does not consider the long-term effects of decisions.
In recent years, RL has become increasingly popular and has also been applied to business process optimization. RL is an optimization method for sequential decision-making problems, such as resource allocation, where a decision-making agent learns a policy by taking actions in an environment sutton_reinforcement_2018 . At each decision step, the agent receives specific information from the business process called the state and, based on the action taken, receives a reward signal. Huang et al. huang_reinforcement_2011 proposed RLRAM, a method that uses the RL algorithm Q-learning, where a Q-table stores the expected future rewards for each state-action pair. They use action-dependent features in their state representation and train their agent to minimize the flow time of the process. This method was extended by Firouzian et al. firouzian_cycle_2019 , who added entropy-based features mined from the event log. While this method works relatively well for problems with small state spaces, the method becomes unsuitable for processes with large state spaces, as separate Q-values are stored for each state. Żibowski et al. zbikowski_deep_2023 solved this issue by DRL to maximize the number of completed cases in a business process. The DRL approach replaces the Q-table with a neural network approximating the value function. However, their method must learn which actions are feasible in each state, which is computationally expensive and may not converge to competitive policies.
Manufacturing processes are closely related to business processes, and numerous methods have been proposed for resource allocation in manufacturing. Many heuristics exist, such as Earliest Due Date or First-In-First-Out sels_comparison_2012 . While these heuristics are robust, explainable, and computationally efficient, they are less effective in situations for which they are not explicitly developed nie_gep-based_2013 . DRL has also been applied in job shop scheduling cunha_deep_2020 . For example, Luo luo_dynamic_2020 trained an agent to select dispatching rules, and Song et al. song_flexible_2023 solved a dynamic assignment problem by employing graph neural networks and DRL to capture complex structural relationships between different resources and jobs. However, these methods often assume the sequence of activities in a case is known, which is not true for business processes.
In this paper, we present two dynamic methods that address the limitations of existing methods for resource allocation in business processes. Our methods are adaptive to the current circumstances of the process and can learn near-optimal policies for realistically sized business processes.
3 Preliminaries
This section introduces definitions and concepts which we will use in this paper. Section 3.1 provides definitions related to resource allocation in business process, and Section 3.2 formalizes the business process model’s execution.
3.1 Business process models
This paper considers business process models describing the activities that can be performed in an organization and the resources that can perform them, including the arrival process of new cases and the processing times of activities (e.g., the business process in Fig. 1). We assume the process model can be simulated or executed according to some execution semantics, which makes the business process observable. Specifically, at any point in time, we can observe the execution state of the business process and events that change the state of the process, such as the completion of activities or the arrival of new cases. We formalize these aspects using the following definitions.
Definition 1 (Event, Trace)
Let $\mathcal{A}$ be the set of activities, $\mathcal{C}$ the set of cases, $\mathcal{R}$ the set of resources, and $\mathcal{T}$ the time domain. An event $e=(a,c,t,r,l)$ is mapped to an activity $a\in\mathcal{A}$, has a unique case identifier $c\in\mathcal{C}$, is executed at timestamp $t\in\mathcal{T}$ by resource $r\in\mathcal{R}$, and has a life cycle stage $l\in\{start,complete\}$ to indicate if the event represents the start or completion of executing an activity. In line with van der Aalst van_der_aalst_process_2016 , we refer to the different elements $a$, $c$, $t$, $r$ and $l$ of an event $e$ as #$\#_{a}(e)$, #$\#_{c}(e)$, #$\#_{t}(e)$, #$\#_{r}(e)$ and #$\#_{l}(e)$. A trace is a finite, non-empty sequence of events that describes everything that happened in a business case. We define a trace as $\sigma=\langle e_{1},e_{2},\dots,e_{n}\rangle$, where $n$ is the length of the trace $|\sigma|$. Events within a trace are sequential and should be non-decreasing in time such that $\forall_{1\leq i,j\leq n}\#_{t}(e_{i})\geq\#_{t}(e_{j})$.
The process model in Fig. 1 consists of two activities, which can be executed once for each case in the process. We distinguish executions of the same activity, both within and without the same case, using the term activity instance.
Definition 2 (Activity instance)
An activity instance is an occurrence of an activity that belongs to a specific case. One case may have multiple instances of the same activity during its execution. We define the set of activity instances as $\mathcal{K}$, and each activity instance $k\in\mathcal{K}$ belongs to a case $c$. We define $\mathcal{K}_{c}$ as the activity instances of case $c\in\mathcal{C}$ and $\mathcal{K}_{a}$ as the activity instances of activity $a\in\mathcal{A}$. An activity instance has a life cycle $p\in\{waiting,processing,complete\}$ to indicate whether the activity instance is waiting in the queue, being processed, or is completed. We refer to the life cycle stage of an activity instance as #$\#_{p}(k)$.
To model the execution state of a business process for the purpose of learning what the best resource allocation is in a particular state, we must be able to observe what activity instances are waiting for processing and what resources are available to be assigned to these activity instances.
Definition 3 (Unassigned Activity Instance)
An unassigned activity instance is an activity waiting in the queue of an activity to be executed by a resource. We define the set of unassigned activity instances as $K=\{k\in\mathcal{K}\mid\#_{p}(k)=waiting\}$. We use subscripts to indicate to what case $K_{c}$ and activity $K_{a}$ an unassigned activity instance belongs.
Definition 4 (Available Resource, Unavailable Resource)
An available resource is currently not busy executing an activity instance and can be assigned. We define the set of available resources as $R^{+}\subseteq\mathcal{R}$ and the set of unavailable resources as $R^{-}\subseteq\mathcal{R}$ such that $R^{+}\cup R^{-}=\mathcal{R}$ and $R^{+}\cap R^{-}=\emptyset$. A resource is removed from $R^{+}$ and added to $R^{-}$ when executing an activity instance. An unavailable resource is moved from $R^{-}$ to $R^{+}$ after it finishes processing the activity instance to which it was assigned.
The business process execution state contains all information about the condition of the process at a given point in time, such as what resources and activity instances are available, as well as what is currently being processed. From a business process execution state, the process may transition into another state after some time $t$ has elapsed as a consequence of an event happening. These events can be the arrival of a new case, the assignment of an activity instance, the completion of an activity instance, or the completion of a case. Events happen according to the process model. An activity can only be executed if assigned to a resource.
Definition 5 (Business Process Execution State)
A business process execution state is the set of active cases $C\subseteq\mathcal{C}$, unassigned activity instances $K\subseteq\mathcal{K}$, available and unavailable resources $R^{+},R^{-}\subseteq\mathcal{R}$, and current assignments $B\subseteq K\times R^{-}$ that exist at a particular moment in time $t$.
3.2 Business process execution
The resource allocation problem we aim to solve in this paper is a dynamic task assignment problem spivey_dynamic_2004 . Given a business process execution state, we aim to assign a resource to an unassigned activity instance to optimize some KPI under specific constraints. We define the specific parts of the optimization problem, which are the decision variables, objective, and constraints, in the context of the dynamic task assignment below.
One explicit constraint in a business process is resource eligibility, which says that not every resource can perform each activity due to, for example, a lack of a specific skill set or level of authorization.
Definition 6 (Resource Eligibility)
The process model defines what resources are eligible to perform what activity. We refer to the set of resources that can execute activity $a\in\mathcal{A}$ as $\mathcal{R}_{a}\subseteq\mathcal{R}$, which implies that $r\in\mathcal{R}_{a}$ can also execute each unassigned activity instance $k\in\mathcal{K}_{a}$.
The process model in Fig. 1 defines that both resources can perform activity RA, while only resource $r_{1}$ can perform activity AA. Therefore, the eligible resource set for activity RA is $\mathcal{R}_{RA}=\{r_{1},r_{2}\}$, and for activity AA is $\mathcal{R}_{AA}=\{r_{1}\}$.
The decision variables are the assignments, which are the allocation of an available resource to an unassigned activity instance. The set of assignments only includes assignments that satisfy the resource eligibility constraint (Def. 6). However, not all assignments are possible in any given business process execution state (Def. 5) because there must be an unassigned activity instance (Def. 3) and an available resource to perform it (Def. 4). Under these constraints, we define the set of possible assignments as a subset of the set of assignments.
Definition 7 (Assignment, Possible Assignment)
An assignment is a tuple $(r,k)$ of a resource $r\in\mathcal{R}$ and an activity instance $k\in\mathcal{K}$. Considering resource eligibility, we define the set of assignments as $\mathcal{D}=\{(r,k)\mid\forall a\in\mathcal{A},r\in\mathcal{R}_{a},k\in\mathcal{K}_{a}\}$. However, not every assignment is possible in all business process execution states, as there must be unassigned activity instances, and resources must be available and eligible to perform them. We define the set of possible assignments in a specific business process execution state as $D=\{(r,k)\mid(r,k)\in\mathcal{D},r\in R^{+},k\in K\}$.
After making an assignment, completing an activity instance, or upon the arrival of a new case, the business process execution state (Def. 5) transitions to a new state.
Definition 8 (State Transition)
A state transition happens each time a decision is made to assign resource $r$ to unassigned activity instance $k$. First, the business process execution state (Def. 5) is adapted by changing the set of unassigned activity instances to $K-\{k\}$, the set of available resources to $R^{+}-\{r\}$, the set of unavailable resources to $R^{-}\cup\{r\}$, and the set of current assignments to $B\cup\{(r,k)\}$. The set of possible assignments, $D$, also changes depending on $R^{+}$ and $K$. The state also changes after an activity instance completes, which releases the resource of assignment $(r,k)$ such that $R^{+}\cup\{r\}$, $R^{-}-\{r\}$ and $B-\{(r,k)\}$. Furthermore, if after completing an activity instance $k\in\mathcal{K}_{c}$, case $c$ is not completed, a new unassigned activity instance $k$ is generated, i.e., $K\cup\{k\}$. Upon a new case arrival, an unassigned activity instance is also generated.
The last part of the optimization problem is the objective, which is to minimize the average cycle time of cases in the process.
Definition 9 (Completed Case, Cycle Time)
For a sequence of events $\sigma=\langle e_{1},e_{2},\dots,e_{n}\rangle$ of a case, the cycle time $c_{CT}$ is the difference between the time of occurrence of the first and the last event, i.e., $c_{CT}=\#_{t}(e_{n})-\#_{t}(e_{1})$. We can determine the cycle time for an ongoing case #$\#_{c}(e_{i})\in C$ or a completed case #$\#_{c}(e_{i})\in\mathcal{C}-C$.
Thus, the objective function can be written as:
$\operatorname*{minimize}\frac{1}{|\mathcal{C}-C|}\sum_{c\in\mathcal{C}-C}c_{CT}$ (1)
The interaction between the environment and the decision-making agent is as follows. In any given business process execution state (Def. 5), the environment checks for any possible assignments (Def. 7). If there are, the agent is asked to make a single assignment, and the environment transitions to the next business process execution state (Def. 5) by removing the respective available resource (Def. 4) and unassigned activity instance (Def. 3). If there are still possible assignments (i.e., $D\neq\emptyset$), the agent is asked for another assignment. When there are no more possible assignments or if the agent decides to postpone its decision, the business process simulation resumes, which modifies the business process execution state according to Def. 8. After the business process execution state changes, the environment checks for possible assignments again.
4 Method
In this section, we present two learning-based methods for resource allocation in business processes to minimize the average cycle time of cases. First, we present a Deep Reinforcement Learning (DRL) method, and second, we present a Score-based Value Function Approximation (SVFA) method.
The main differences between the two methods are how each algorithm is trained and how information is represented in each algorithm. The DRL method learns continuously during the simulation of the business process based on a reward it receives after each decision step, whereas SVFA receives a single reward signal at the end of the simulation. SVFA’s advantage is that it directly minimizes the objective function. Conversely, rewarding the agent to minimize the objective function in our DRL approach can introduce bias. For example, our reward function might prioritize the quick completion of one case while delaying others, leading to high rewards but detrimental effects on the average cycle time. The information provided in the SVFA method is represented in a set of hand-crafted features based on domain knowledge and optimizes these weights. On the other hand, the DRL agent uses a neural network that automatically learns important features from the state space. The hand-crafted features contain important information for resource allocation but may overlook vital information hidden in the data, which a neural network can discover.
We provide two methods to demonstrate different ways to approach resource allocation for business processes using learning-based methods. Section 4.1 details the DRL method and Section 4.2 the SVFA method.
4.1 Deep Reinforcement Learning
According to Sutton & Barto sutton_reinforcement_2018 , RL is a method to train one or multiple agents to interact with an environment to maximize a reward signal. The term DRL is used for RL algorithms that employ neural networks as function approximations. Fig. 2 shows how we developed the reinforcement cycle for resource allocation in business processes. The environment is a discrete-event simulation of the business process. Each time an assignment is possible, according to Def. 7, the simulation prompts the agent to make a decision, which is an assignment. We refer to this moment as a decision step. The environment processes the agent’s chosen action, and based on the action’s effect on the environment, a reward is returned to the agent. One run of the simulation is called an episode and can contain a different number of decision steps, depending on the stochastic events that take place in each run. The agent learns a policy, which is a function that maps the state to the actions and defines which action should be taken in a given state to maximize the reward.
Figure 2: Agent’s perspective of the agent-environment interaction, adapted from sutton_reinforcement_2018 .
In this paper, we use the state-of-the-art DRL algorithm Proximal Policy Optimization (PPO) schulman_proximal_2017 . Depending on the available resources and unassigned activity instances, several assignments may not be possible at a specific decision step (Def. 7). Therefore, we use a masking function that, given a state, prevents the agent from taking these infeasible actions, improving training speed and scalability huang_closer_2022 . We model the resource allocation problem as a finite Markov Decision Process (MDP) within the DRL framework. The remainder of this section defines each component of the MDP, as shown in Fig 2.
State
The business process execution state (Def. 5) contains all information related to the current condition of the process. The state, as defined in the RL paradigm, describes what the agent observes from the business process execution state. To create the features in the state, we order the resources $r\in\mathcal{R}$ and activities $a\in\mathbf{A}$. We use subscript $r_{i}$ and $a_{j}$ to indicate the position of a specific resource and activity in the list, where $i\in\{1,\dots,|\mathcal{R}|\}$ and $j\in\{1,\dots,|\mathcal{A}|\}$. The state space contains:
    A mapping $\delta_{i}\rightarrow\{0,1\}$, indicating for each resource whether it is assigned or not (Def. 4), e.g., for resource $r_{i}:1$ if $r_{i}\in R^{+}$, or 00 if $r_{i}\in R^{-}$.
    A mapping $\eta_{i}\rightarrow[0,1]$, indicating for each resource which activity the resource is assigned to (Def. 7), e.g., for resource $r_{i}:\{\frac{j}{|\mathcal{A}|}\mid k\in K_{a_{j}}\}$, if $(r_{i},k)\in B$, or 0 otherwise.
    A mapping $\kappa_{j}\rightarrow[0,1]$, indicating the queue of unassigned activity instances of each activity. To reduce the dimensionality of the state space, the queue is normalized by dividing the number of unassigned activity instances by 100 and truncating values higher than 1, i.e., for activity $a:\min(\frac{|K_{a}|}{100},1)$.
The input of the agent is the state vector. As shown above, the features in the vector are normalized between 0 and 1, which is a common practice to improve sample efficiency during training henderson_deep_2018 . Given a state, the agent receives information about which resources are available, how many unassigned activity instances are queued at each activity, and the current assignments. Using this information, the agent can learn what assignments lead to high rewards based on the state. Furthermore, it can learn when other resources are likely to be available again, and it can, for example, postpone a decision to wait for a better assignment. The size of the state vector is $2\cdot|\mathcal{R}|+|\mathcal{A}|$.
Action
The actions are all assignments (Def. 7) and a postpone action, which delays the current decision step and evolves the environment to the next decision step. At each decision step, a single action is taken. As mentioned, we mask infeasible actions ($\mathcal{D}-D$ according to Def. 7). The masking function is a vector that maps $\mathcal{D}\rightarrow\{0,1\}$, indicating if an assignment is possible or not, i.e., $1$ if $(r,k)\in D$ and 00 otherwise. The postpone action is always feasible. The action vector is multiplied by the masking vector, setting the probability of choosing infeasible actions to zero. The size of the action vector is $|\mathcal{D}|+1$.
State transition
Given a specific state and action, the business process transitions to a new state, according to Def. 8. Multiple actions can be taken at the same moment in time, depending on the availability of resources or unassigned activity instances. In this case, actions are taken one at a time until no assignments are possible or the agent chooses the postpone action. If there are no possible assignments, the environment evolves until the next decision step. If the agent postpones, the environment evolves until there is at least one different unassigned activity instance or available resource compared to the previous state.
Reward
The optimization objective is to minimize the average cycle time of cases (Def. 9), and we should reward the agent for actions that contribute to this objective. Therefore, we give a reward of $\frac{1}{c_{CT}+1}$ for each case $c$ that is completed. We inverse the cycle time such that a low cycle time results in a high reward and add $+1$ to the denominator to reduce the dimensionality of the rewards.
Having defined these elements, we can train an agent that learns a policy function for resource allocation that minimizes the average cycle time of cases. The implementation details of the algorithm are given in Section 5.2.1.
4.2 Score-based value function approximation
The second method we propose is a score-based value function approximation (SVFA) approach that, based on a set of features, learns to minimize the average cycle time of cases. We consider all possible assignments $(r,k)\in D$ (Def. 7) in a given business process execution state and compute a score for each of them. The score is a function of a set of features related to assignment $(r,k)$ and learned weights for each feature. The features are hand-crafted based on expert knowledge. The features are:
    $MeanAssignment(r,k)$: the mean processing time of the assignment.
    $VarAssignment(r,k)$: the variance of the processing time of the assignment.
    $ActivityRank(r,k)$: we evaluate $MeanAssignment(r,k^{\prime})$ of resource $r$ for all other unassigned activity instances $k^{\prime}\in K$ and rank them on the mean processing time, where the lowest mean processing time is ranked the highest. $ActivityRank(r,k)\rightarrow\mathbb{N}$ is the rank of unassigned activity instance $k$.
    $ResourceRank(r,k)$: we evaluate $MeanAssignment(r^{\prime},k)$, for all other available resources $r^{\prime}\in R^{+}$ and rank them on the mean processing time, where the lowest mean processing time is ranked the highest. $ResourceRank(r,k)\rightarrow\mathbb{N}$ is the rank of resource $r$.
    $ProbFin(r,k)$: the probability that after an assignment $(r,k)$ case $c$ is completed, where $k\in K_{c}$.
    $QueueLength(k)$, the number of unassigned activity instances for a specific activity $|K_{a}|$, given that $k\in K_{a}$.
In this paper, features such as $MeanAssignment(r,k)$, $VarAssignment(r,k)$, and $ProbFin(r,k)$ are determined by the business process model. However, an estimate may also be derived from historical data if available. The score function maps each state-action pair to a real value, which represents the goodness of making a specific assignment in a specific state:
$Score(r,k)=w_{1}\cdot MeanAssignment(r,k)$ (2)
$+w_{2}\cdot VarAssignment(r,k)+w_{3}\cdot ActivityRank(r,k)$
$+w_{4}\cdot ResourceRank(r,k)-w_{5}\cdot ProbFin(r,k)$
$-w_{6}\cdot QueueLength(k),\textrm{where }w_{i}\geq 0,\textrm{for}i\in\{1,\dots,6\}$
To calculate the score, we aggregate the first four features in Equation 2 through addition because a lower value is better (i.e., lower cycle time). A high value for the last two features $ProbFin(r,k)$ and $QueueLength(k)$ is preferred, as finishing cases and minimizing queue lengths result in a lower average cycle time hopp2011factory . Therefore, these features are subtracted when calculating the score.
At each decision moment, the score for all possible assignments (Def. 7 is calculated. The assignment with the lowest score is the best and will be made unless its score does not exceed a learned threshold $w_{7}$. Instead, the agent postpones if no assignment’s score is lower than this seventh weight. Similarly to the DRL approach, when multiple assignments are possible, the agent makes assignments one at a time and then reevaluates all scores based on the new state to make the next assignment.
We use Bayesian optimization to find weights that minimize the average cycle time. Bayesian optimization is a global optimization technique that uses a probabilistic model to predict the relationship between the input parameters and the output of a function. The algorithm learns as new data is obtained and focuses on selecting parameters in promising areas of the search space. For more information on the Bayesian optimization algorithm, we refer to Frazier frazier2018tutorial . The implementation details are given in Section 5.2.2.
5 Evaluation
Section 5.1 introduces six business processes with archetypal process flows and characteristics and three realistically sized business processes on which we evaluate our methods. Section 5.2 describes the implementation details of our methods and the hyperparameters used for training them. Section 5.3 details the evaluation protocol and introduces the benchmarks. Section 5.4 shows the results of our methods compared with three heuristic benchmarks and two benchmarks from the literature.
5.1 Experimental setup
This section presents the business processes on which we evaluate and benchmark our methods. First, we design six different business processes, each consisting of two activities and two resources, to analyze which methods work well in what specific situations. We refer to these six business processes as scenarios. Each scenario is characterized by process patterns such as parallel or choice constructs, varying utilization rates, and resource competition. Second, we connect these six scenarios sequentially and in parallel to create three realistically sized business processes containing all the different process flows seen in the scenarios. Evaluation of these business processes demonstrates which methods perform well under specific circumstances (i.e., the scenarios) and which methods perform well when all scenario dynamics are combined. We refer to these three business processes as composite business processes.
We first introduce the six scenarios. In each scenario, cases arrive according to a Poisson process with rate $\lambda=\frac{1}{2}$ time units. The processing times are exponentially distributed with mean $X_{r}$, $r\in\mathcal{R}$. The mean processing times are chosen to align with the characteristics of their respective scenario (e.g., utilization level, a slower resource) and ensure that each scenario’s cycle time is in the same order of magnitude. The latter is important for connecting the scenarios to create the composite business processes such that a single scenario does not dominate the total cycle time of the composite process. The six scenarios are shown in Fig. 3.
(a) Scenario 1: Low utilization
(b) Scenario 2: High utilization
(c) Scenario 3: Slow server
(d) Scenario 4: Slow downstream
(e) Scenario 5: N-network
(f) Scenario 6: Parallelism
Figure 3: Six business process models with different characteristics used for evaluation.
Fig. 2(a) and Fig. 2(b) show the first two scenarios characterized by their utilization rates. The process’s utilization rate significantly impacts the average queue length at an activity hopp2011factory , i.e., $|K_{a}|$. If the utilization rate is low, such as in scenario 1 (Fig. 2(a)), the average size of the queues in the process will also be smaller, and therefore, the agent should focus on minimizing the processing time. Scenario 2 in Fig. 2(b) has a high utilization rate, which means that the average queue length will be higher, and cases will spend most of their cycle time queuing. Therefore, minimizing queueing time becomes more important in this scenario.
Fig. 2(c) shows scenario 3 in which an experienced employee ($r_{5}$) works faster than an inexperienced employee ($r_{6}$). An efficient assignment policy should avoid assigning the inexperienced to activity $F$, as this resource is especially slow at executing it. Scenario 4 in Fig. 2(d) shows a business process where both resources have the same processing time distribution for each activity, of which the second activity, $H$, is processed slower than the first. Therefore, the prioritization of activities is more important than the resource that executes them.
Fig. 2(e) models scenario 5, a process with two activities with different resource eligibility. This process is a variant of the well-studied bipartite matching problem (e.g., Cadas et al. cadas_bipartite_2019 ). It is also known as the ‘N-network’, referring to the ‘N’ shape of the bipartite graph, where nodes represent resources and activities, and edges represent the resource eligibility (Def. 6). The decision XOR gateway characterizes the process, where cases go to either activity with probability $p=0.5$. This scenario represents a situation where one employee ($r_{10}$) can perform more activities due to, for example, the authorization level. This resource’s capacity should be balanced over the activities. Fig. 2(f) shows scenario 6, which contains a parallel process flow. In this scenario, both activities can be performed simultaneously, but a case is only completed once both activities have been executed. The processing time at each activity is the same for both resources, making prioritizing what activity to complete the most important choice.
In the second part of the evaluation, we create three realistically sized business processes by combining the six scenarios. We refer to these as composite business processes. The first composite business process is created by sequentially connecting the scenarios, starting with scenario 1 and ending with scenario 6. In the second composite business process, we also connect the scenarios sequentially but in reverse (i.e., starting with scenario 6 and ending with scenario 1) to investigate the impact of the order of the scenarios. Note that the external arrival process is still a Poisson process, while the internal arrivals now depend on departures from previous activities. The third composite business process is created by putting all scenarios in parallel.
5.2 Training setup
In this section, we describe the training setup that we used to train our methods. Section 5.2.1 describes the training procedure and the hyperparameters used for the DRL approach. Section 5.2.2 describes how SVFA’s optimal feature weights are obtained for each business process.
5.2.1 Deep Reinforcement Learning
We performed hyperparameter tuning for the sequential composite business process using grid search and used the found optimal hyperparameters to train an agent for all business processes considered in this paper. The hyperparameter ranges and optimal values are shown in Table 1.
Table 1: Hyperparameter ranges for grid search on the composite business process and optimal hyperparameters found for training maskable PPO agent.
Parameter Range Optimal
Number of layers [1, 2] 2
Neurons per layer [32, 64, 128] 128
Clip range [0.1, 0.2] 0.2
Number of samples per update [25600, 51200] 25600
Batch size [256, 512] 256
Learning rate [3e-3, 3e-4, 3e-5] 3e-5
Discount factor [0.99, 0.999, 0.9999] 0.999
After tuning the hyperparameters, we found that a high number of observations per update and a large batch size significantly improved the model’s performance. We set the number of samples before updating the model to include at least an entire episode. In the composite business processes, there are 11 activities per case, requiring at least the same number of actions before the case is completed and a non-zero reward is returned. When using a larger batch size, more of the delayed and sparse rewards are included in each model update, which gives a more accurate estimate of the gradient and helps reduce the sample noise.
During training, we decrease the learning rate linearly with the total training steps to encourage exploration at the start and exploitation of the policy at the end of training. The discount factor determines the importance of future rewards. We found a high value to be optimal to account for the long-term effects of actions on the cycle time and sparse rewards. We evaluated the current model during training and saved the best model found.
In the N-network, the DRL algorithm converged to a policy where the agent only postpones its decisions. Each case in this scenario only contains a single activity before the case is completed, and the agent receives a reward. Because of this, the agent learned to prioritize one of the two activities and get high rewards by completing that activity quickly. To accomplish this, the agent learned to postpone instead of allocating a resource for the other activity. While this policy may result in high rewards, it is obviously not beneficial for the mean cycle time. To mitigate the bias induced by the reward function, we impose a penalty of -0.1 every time the agent uses the postpone action in this scenario. Penalizing the agent in this scenario led to competitive results with the best-performing benchmarks, which will be shown in Section 5.4.
We use the maskable PPO implementation of the contributory version of the popular DRL Python library Stable-Baselines3  [^1] https://github.com/Stable-Baselines-Team/stable-baselines3-contrib and create our environment using Gymnasium  [^2] https://gymnasium.farama.org/index.html. Untuned hyperparameters were kept at their default values.
5.2.2 Score-based value function approximation
To train the weights of the SVFA method, we use Bayesian optimization. We define a search space in which we optimize the value function’s weights for each business process. In our setup, we use $0\leq w_{i}\leq 100$, for $i\in\{1,\dots,7\}$. The optimization procedure runs for a maximum of 1000 trials and converges before using all trials in all business processes. The method was implemented in Python using the scikit-optimize package [^3] https://scikit-optimize.github.io/stable/.
5.3 Evaluation protocol and benchmarks
This section describes our approach to evaluate each method and the benchmarks. We first describe how we evaluated the methods and then introduce the benchmarks and their implementation specifics.
Each method is evaluated on 100 simulations of 5000 time units in each business process. The cycle time is calculated according to Equation 1. Uncompleted cases at the end of the process are truncated and included in the average cycle time calculation. This truncation ensures that the best policy is not always to postpone, as otherwise, no case completions would result in a zero average cycle time. We test if the method with the lowest cycle time significantly differs from the other methods using a T-test for two independent samples ($p\geq 0.05$).
To benchmark our methods, we compare them with three traditional heuristics and two methods from the literature park_prediction-based_2019 ; zbikowski_deep_2023 . Not all benchmarks were directly implementable for our problem, or no source code was available, therefore, we had to make small changes from the original implementation. The benchmarks and their implementation details are:
    Shortest Processing Time (SPT): we evaluate all possible assignments at a decision step on their expected processing time and choose the assignment(s) with the shortest processing time.
    First In, First Out (FIFO): the case that is in the business process for the longest amount of time is prioritized at every activity.
    Random: a random assignment is made from all possible assignments at a decision step. This heuristic serves as a baseline.
    Park and Song park_prediction-based_2019 : this method uses a prediction model to predict the next activity and the processing time of possible assignments. Since we do not have an event log to train a prediction model, we derive the next activity and the expected processing time of each assignment from the business process model. Therefore, the predictions we passed to the scheduling stage are 100% accurate.
[…]
We chose these benchmarks because their methods are closely related to our methods and are suitable for the online scheduling problem. Considering we do not need to make any predictions, the minimum-cost maximum-flow algorithm by Park and Song park_prediction-based_2019 is an advanced version of SPT and, therefore, a relevant benchmark. We implemented RLRAM huang_reinforcement_2011 as a preliminary benchmark. However, the method does not scale well to business processes containing much stochasticity like those considered in this paper. Our implementation of RLRAM failed to converge on competitive policies within a reasonable time frame. Our implementation of RLRAM algorithm can be found in our code. Other methods from the literature, mine features from an event log to solve the resource allocation problem firouzian_cycle_2019 ; xu_resource_2008 ; zhao_optimization_2015 ; arias_framework_2016 ; kuchar_automatic_2016 . However, our evaluation framework uses simulated data without a real-world event log; therefore, these event log-based approaches could not be directly replicated.
5.4 Results
Table 2: Mean cycle time and half-width of the 95%-confidence interval for the six scenarios for DRL and SVFA, compared with three heuristics and two existing methods. Results were obtained over 100 simulations. A lower cycle time is better. The best result is in bold, and multiple bold results indicate they are not significantly different from the best method.
Business process DRL SVFA SPT FIFO Random Park & Song park_prediction-based_2019 Żbikowski et al. zbikowski_deep_2023
Low utilization 5.7 (0.09) 5.6 (0.08) $5.8\,(0.08)$ $5.9\,(0.11)$ $6.6\,(0.14)$ $6.0\,(0.09)$*
High utilization 18.9 (0.91) 18.1 (0.94) 18.7 (0.89) $27.0\,(2.77)$ $35.8\,(4.20)$ 19.4 (1.01) -*
Slow server 12.2 (0.38) 11.4 (0.31) $27.1\,(1.50)$ $19.3\,(1.33)$ $21.9\,(1.49)$ $20.5\,(1.26)$ $58.9\,(5.60)$
Slow downstream 9.9 (0.33) 9.6 (0.25) $15.4\,(0.71)$ 9.8 (0.35) $11.9\,(0.53)$ $14.7\,(0.59)$*
N-network $5.9\,(0.11)$ 5.7 (0.09) $7.4\,(0.20)$ $5.9\,(0.11)$ $6.5\,(0.16)$ $6.3\,(0.14)$ $6.7\,(0.11)$
Parallelism $11.3\,(0.38)$ $12.0\,(0.38)$ $13.8\,(0.57)$ 9.3 (0.31) $11.0\,(0.45)$ $13.7\,(0.60)$*
    Learned policy only selected infeasible actions, resulting in indefinite postponement.
As can be seen in Table 2, the DDQN method by Żbikowski et al. zbikowski_deep_2023 could not converge to a competitive policy in some scenarios. The reason for this is that this algorithm can get stuck in a loop where it only selects infeasible actions. In some cases, the representation of their state space does not change until an assignment is made (see zbikowski_deep_2023 ). If the deterministic policy takes an infeasible action in that state, no assignments are made, and thus, the state never changes, resulting in indefinite postponement.
Table 2 shows that both of our methods outperform or are competitive across benchmarks, except for the scenario involving parallelism. In the low utilization scenario, strategic decisions like idling or prioritizing case completions become more viable as the load on the resources is. Our methods excel in learning these strategies. On the other hand, in the high utilization scenario, prioritizing the assignment of activity instances to the fastest available resources is crucial to prevent excessive queue buildup. This strategy aligns closely with the SPT heuristic and with Park and Song’s method park_prediction-based_2019 , as shown in these methods’ performance. Our methods can also learn this policy effectively.
In the slow server scenario, resource $r_{5}$ is inefficient in executing activity E and avoiding this assignment whenever possible is advisable (see Fig. 2(c)). Hence, strategically idling to await a faster resource’s availability may be beneficial. Park and Song’s method park_prediction-based_2019 also accommodates strategic idling, which is evident in its improved performance over SPT. Both of our methods successfully learned policies that outperform all benchmarks.
In the slow downstream scenario, the processing time of an activity is equal across resources. Therefore, the emphasis shifts from choosing what resource to allocate to what activity to execute. Finishing cases as quickly as possible leads to the lowest cycle time, as seen in the FIFO heuristic’s high performance. Our methods are also capable of learning this policy.
In the N-network, only resource $r_{10}$ can execute activity I (see Fig. 2(e)); therefore, reserving it for that activity is an effective strategy. While SVFA was able to find a policy that outperforms the other methods, our DRL model struggled. The DRL agent learned only to execute activity J and postpone otherwise, leading to high immediate rewards. However, this policy resulted in a large queue building up at activity I, and, therefore, a high average cycle time. After introducing a penalty of $-0.1$ for postponing in this scenario, the agent converged to a better policy.
In the parallelism scenario, a FIFO policy yielded the best performance. FIFO completes all activities of a case before initiating the next case, resulting in less waiting time and lower cycle times. In contrast, other methods lack information about what activity instances lead to case completions, making finding an optimal resource allocation policy more challenging. The results for the composite business processes are shown in Table 3.
Table 3: Mean cycle time and half-width of the 95%-confidence interval for composite business processes for DRL and SVFA, compared with three heuristics and two existing methods. Results were obtained over 100 simulations. A lower cycle time is better. The best result is in bold, and multiple bold results indicate they are not significantly different from the best method.
Business process DRL SVFA SPT FIFO Random
Composite 64.7 (2.63) $72.1\,(2.34)$ $98.8\,(3.80)$ $69.0\,(3.36)$ $84.9\,(4.00)$
Composite reversed 62.3 (1.77) $73.6\,(2.44)$ $106.6\,(5.18)$ $68.1\,(2.61)$ $85.0\,(4.74)$
Composite parallel 23.4 (0.92) $34.1\,(1.29)$ $35.1\,(1.73)$ $31.0\,(2.54)$ $37.1\,(2.6)$
Similarly, as seen in some of the scenarios (Table 2), the method by Żbikowski et al. zbikowski_deep_2023 converged to a policy that postponed indefinitely. Furthermore, the inference time of the method by Park and Song park_prediction-based_2019 was intractable; therefore, we could not evaluate the method for the composite business processes. This limitation was also indicated by the authors. For this reason, the results of these two benchmarks are not included in the evaluation and Table 3.
Table 3 shows that DRL outperforms all methods in the composite business processes. The performances of the composite and composite reverse business processes are similar, which means that the order of scenarios does not significantly impact the performance. This was to be expected because the scenarios in the composite business processes are still somewhat isolated, as each scenario has its own resources. While FIFO outperformed the other methods in the parallel scenario, as seen in Table 2, it could not effectively handle the dynamics of all scenarios combined, as seen in Table 3.
SVFA performed well in the scenarios but could not find competitive policies in the composite business processes. A possible explanation is that there are competing feature weights for each scenario. For example, as can be seen in Table 2, an SPT policy yields good performance in the high utilization scenario, from which we conclude that the weight of the feature $MeanAssignment$ should be high for this scenario. On the other hand, SPT makes poor assignments in the slow downstream scenario, which means the weight of $MeanAssignment$ should be low in this scenario. A balance must be found in the composite business processes for this feature to account for both scenarios, which can result in a sub-optimal policy.
6 Conclusion
This paper presented two learning methods for resource allocation in business processes. The first approach is based on deep reinforcement learning (DRL), where an agent learns what assignments lead to a low cycle time by interacting with the environment. The second method is a score-based value function approximation (SVFA) approach, where a score function is learned that ranks assignments based on a set of curated features derived from the current process circumstances. We evaluate our proposed methods by benchmarking them against three traditional heuristics and the two most relevant existing BPO methods. We evaluate the methods on six business processes, also known as scenarios, representing archetypal process patterns and three realistically sized business processes that are a combination of these six scenarios.
The evaluation results show that our learning methods outperform or are competitive with the benchmarks in five out of six scenarios. Only in the scenario with parallelism did the First In, First Out heuristic outperform our methods because it prioritizes the activity that leads to case completion. Our methods and other benchmark methods do not explicitly consider information related to the execution sequence of cases, resulting in lower performance. One of the benchmarks from the literature was competitive with our methods in the scenario with high utilization, but neither of these existing methods could find a competitive policy for the other scenarios. These results showed that our learning-based methods are suitable for learning policies for process patterns commonly seen in business processes.
The second part of the evaluation was performed on business processes created by combining the six scenarios in three different configurations: in sequence, in sequence with reversed order, and in parallel. While some heuristics were competitive or outperformed our methods in the scenarios, the heuristics exhibited limited performance when all process patterns from the scenarios were combined. Our proposed DRL method outperformed all methods in all three composite business processes and outperformed the best-performing benchmark by, on average, 13.1%. DRL automatically learns important features from the current state of the process, independent of the process characteristics, making it adaptive to different business process patterns and flows. Our DRL-based method is scalable and suitable for learning resource allocation policies in business processes. While SVFA was, on average, the best-performing method in the scenarios, it could not find a competitive policy in the composite business processes compared with DRL.
In future work, the current methods can be improved in several ways. One research direction is to consider other or multiple optimization objectives. While this work only focused on minimizing the cycle time, other objectives can be a more fitting choice depending on the setting. Furthermore, we trained and tested each model in the same environment. However, in a real-world process, resources may, for example, enter and leave the process, which current methods do not account for. Therefore, handling changes in the business process model is a relevant extension of our work. The process mining field has shown that the execution sequence of a running case, or prefix, holds valuable information but has not been used in any BPO approach so far. The scenario with the parallelism structure showed that the prefix can significantly improve performance, and we will investigate how we can use the prefix in our resource allocation algorithms. Furthermore, one limitation of DRL is that it is difficult for the agent to learn to what extent what action contributes to a specific reward realized many actions later. This behavior is caused by the design of the reward function. While it worked well in learning policies for the business processes considered in this paper, it may also lead to undesired behavior, as seen in the N-network. In future work, we aim to address the bias induced by the reward function.
Reproducibility. The source code to reproduce the experiments can be found in our Github repository.
Acknowledgements. The research leading up to this paper is supported by the Dutch foundation for scientific research (NWO) under the CERTIF-AI project (grant nr. 17998).